# Invariance transformations in wave-optics lensing: implications for gravitational-wave astrophysics and cosmology

### v1

@@ I Introduction
Gravitational lensing of gravitational waves (GWs) occurs when the signal passes close to a massive object that deflects it. This effect could result in multiple images of the same event, magnifications, interference patterns, and delays w.r.t. the unlensed signal. With the advancement of the fourth observing run (O4) of the network of LIGO (with the Livingston and Hanford detectors Aasi et al. (2015)), Virgo Acernese et al. (2014), and KAGRA Akutsu et al. (2020); Aso et al. (2013); Somiya and (for the KAGRA Collaboration) (LVK) and in future runs Abbott et al. (2016), the number of GW detections will grow further, and with it the probability of detecting a lensed event. For this reason, in upcoming observing runs, it will be crucial to be able to account for the potential effects of lensing on observed signals. Neglecting it could introduce biases into our findings regarding the properties of sources Cao et al. (2014); Kim et al. (2024); Shan et al. (2024), as lensing magnification, for example, alters the inferred luminosity distance. Furthermore, under strong lensing conditions, distortions in the lensed signal may occur, particularly when higher modes, precession, or eccentricity are significant factors Ezquiaga et al. (2021). In the wave-optics regime (i.e., when the signal size is comparable to the size of the lens), the modulation of the signal could mimic some feature of the source, like, for example, the eccentricity, as well as biasing other parameters Mishra et al. (2024). To ensure accurate analyses, it will be necessary to incorporate lensed waveforms into the waveform models for parameter estimation Nakamura (1998); Hannuksela et al. (2019); Lai et al. (2018); Pang et al. (2020); Dai et al. (2020). Alternatively, employing different methods such as Bayesian analysis Liu et al. (2021a); Lo and Magaña Hernandez (2021); Janquart et al. (2021) or deep learning approaches Levasseur et al. (2017); Singh et al. (2018); Kim et al. (2020) may prove beneficial. Lensing effects also warrant consideration in population studies, as they influence inferred redshift and mass distributions by amplifying signals and allowing the detection of events from greater distances Oguri (2018); Xu et al. (2022).
@@
The first observation of a lensed GW is still elusive after searches in the O1 and O2 runs Hannuksela et al. (2019); McIsaac et al. (2020); Li et al. (2023); Dai et al. (2020); Liu et al. (2021b), O3a Abbott et al. (2021), the full O3 run Abbott et al. (2023a), and an extended O3 follow-up analysis Janquart et al. (2023) yielded no confident signatures. Nonetheless, there are good reasons to be optimistic about future detections, as many different studies forecast $\sim 1$ strong lensing detection per year Li et al. (2018); Ng et al. (2018); Xu et al. (2022); Wierda et al. (2021).
@@
According to Bondarescu et al. (2023), the probability of microlensing (i.e. in the wave-optics regime) is $<20\%$ of all detectable events above the threshold signal-to-noise ratio (SNR)111 In the paper, the threshold SNR is set at $\rho=10$. The level of the detectability of a lensed event is influenced by the mismatch between the lensed signal and the (best matching) unlensed waveform because a mismatch $>0$ decreases the detected SNR. for $M_{det}=120\leavevmode\nobreak\ M_{\odot}$ and $<5\%$ for more common events with $M_{det}=60\leavevmode\nobreak\ M_{\odot}$ (where $M_{det}$ is the detector-frame mass).
@@
The detection and subsequent study of lensed GWs is a powerful tool for both astrophysical and cosmological studies. For example, gravitationally lensed events can be useful because (i) contrarily to unlensed GWs, they give an improved localisation of the source Hannuksela et al. (2020); Yu et al. (2020); (ii) the magnification due to the lensing effect extends the range of detectability of GWs to higher redshift (due to magnification) (see e.g. Xu et al. (2022)); (iii) through the characterisation of the lens, one can study the content and composition of the universe Lai et al. (2018); Diego et al. (2019); Oguri and Takahashi (2020); Cremonese et al. (2023); the study of one or several lensed systems can be exploited for (iv) precision cosmology Sereno et al. (2011); Liao et al. (2017); Cao et al. (2019); Li et al. (2023); Hannuksela et al. (2020), (v) statistical cosmology Xu et al. (2022), and (vi) tests of general relativity Baker and Trodden (2017); Collett and Bacon (2017); Fan et al. (2017); Goyal et al. (2021); Ezquiaga et al. (2021); Goyal et al. (2023).
@@
The time delay between multiple images in gravitational lensing has been a long promise to measure the Hubble constant $H_{0}$ Refsdal (1964a), since the time-delay distance, which is a combination of the angular diameter distances between source, lens, and observer, is proportional to it. This method, also called the time-delay cosmography, was first proposed by Refsdal in 1964 Refsdal (1964b) and can be similarly performed with strongly-lensed GWs. For multiple lensed GW signals from compact binary coalescence, the lens model can be inferred from the magnification ratio and the arrival time delay between them, using ground-based detectors Liao et al. (2017); Hannuksela et al. (2020); Wempe et al. (2022), space-based antennas Sereno et al. (2011), or pulsar timing arrays Cremonese and Salzano (2020). If accompanied by electromagnetic (EM) counterparts, the host galaxies of LVK GW sources can be pinpointed using the information from galaxy catalogs Liao et al. (2017); Hannuksela et al. (2020), so that $H_{0}$ could be measured at high accuracy.
@@
Unfortunately, the usefulness of studying lensed events has a large caveat that goes under the name of mass-sheet degeneracy (MSD) Falco et al. (1985); Gorenstein et al. (1988); Schneider and Sluse (2013, 2014). The MSD relies on the fact that the simultaneous scaling of the lens mass and the source plane leaves the geometrical lensing outputs unchanged. Thus, in practice, we have an intrinsic indeterminacy when we analyze observational data to reconstruct the mass (model) of the lens from the image structure. This is a well-known problem in electromagnetic (EM) lensing.
@@
In EM studies, the MSD is always defined in the geometrical optics approximation, when the wavelength of the lensed signal is much smaller than the typical size of the lens. In this limit, the MSD can be broken, or at least attenuated, when: (i) the same gravitational object act as lens for multiple sources at different redshifts Saha et al. (2000); Lubini et al. (2014); Grillo et al. (2020); Liao et al. (2017); (ii) independent mass estimations of the lens are available, e.g. from spatially resolved kinematic observations of the lens Schneider and Sluse (2013); Birrer et al. (2020); Ding et al. (2021); (iii) by a proper combination of the previous astrophysical data with cosmological geometrical data Chen et al. (2020).
@@
In the context of gravitational lensing of GWs, given the much lower frequency of the signal, the wave optics regime can also be relevant as GW wavelengths can be comparable to the lens size Takahashi (2017); Cremonese and Mörtsell (2018); Cremonese and Salzano (2020); Jung and Shin (2019); Dai et al. (2018); Sun and Fan (2019); Liao et al. (2020). Even though the MSD can be broken in this limit Cremonese et al. (2021), such a degeneracy can affect lens mass modeling and can highly bias the estimation of the Hubble constant, $H_{0}$ Schneider et al. (1992); Schneider and Sluse (2013); Kochanek (2020), contributing to the systematic error budget. This is the reason why the MSD is carefully analyzed in projects like H0LiCOW Chen et al. (2019); Wong et al. (2020), TDCOSMO Millon et al. (2020) and in the Time Delay Lens modeling Challenge Ding et al. (2018, 2020). In GW lensing, this topic was first discussed in Cremonese et al. (2021) and, more recently in the geometric optics regime in Poon et al. (2024).
@@ II.1 General phenomenology and formalism
Therefore, under these approximations, the problem simplifies to that of solving a scalar wave equation on a medium, i.e. a Helmholtz equation (see e.g. Nakamura and Deguchi (1999) for a review):
@@
Apart from the phase shift of type II images that can induce waveform distortions on signals with multiple frequency components Ezquiaga et al. (2021), lensing is achromatic in the SPA.
@@
In other words, weak-field lensing occurs at scales much larger than the strong gravity regime unless the source is very close to the lens, $D_{LS}\ll D_{L},D_{S}$, as is the case e.g. for the photons near the super-massive black holes imaged by the Event Horizon Telescope Akiyama et al. (2019, 2022). For typical astrophysical lensing situations in which $D_{LS}\approx D_{L}\approx D_{S}$, then
@@ II.2 GW lensing
  • the signal is understood from first principles: general relativity can be used to derive waveform models for CBCs (though numerical simulations and various effective modeling approaches are needed in practice, see Schmidt (2020)),
@@
All these characteristics will be relevant for different aspects of gravitational lensing. Importantly, GW sky localizations, which are at best of the order of few square degrees in optimistic current observing scenarios Abbott et al. (2016), will be much larger than the typical angular scale of the problem $\theta_{*}$.
@@ III Invariance Transformations in gravitational wave lensing
Our ability to learn about astrophysics and cosmology from lensed systems depends on how well we can distinguish the predictions of different models on the observables. For this reason, we now focus on studying which sets of variations of the lens system leave the observables unchanged, as those will lead to degeneracies between different astrophysical and cosmological interpretations. Such “invariance transformations” have been widely studied in the EM spectrum since they constitute a large fraction of the error budget when reconstructing the lens model or inferring the expansion rate of the Universe Falco et al. (1985); Gorenstein et al. (1988); Schneider and Sluse (2013, 2014). They have also been explored in the context of GWs Cremonese et al. (2021); Poon et al. (2024), and here we will build upon those results.
@@ III.1 Multiple-image regime
Following Schneider and Sluse (2013), we are going to consider three possible modifications of our fiducial lens scenario:
@@ III.2 Wave-optics regime
We now explore these invariance transformations in the regime of wave optics. Since the diffraction integral is not analytically solvable in general, we restrict to the simpler case of spherically symmetric lenses. In this case, the diffraction integral in Eq. (7) can be simplified to one dimension, integrating over the angular polar coordinate. Doing so, one arrives at Nakamura and Deguchi (1999)
@@ III.3 Point mass model
To gain further intuition, we fully specify the lens model to exemplify the effect of the invariance transformations. We start with the simplest model: a point mass (PM), where the potential is given by $\Psi(x)=\log x$. The Einstein radius in the PM model is $\theta_{*}=\sqrt{4GM_{l,z}/c^{3}\tau_{D}}$ Takahashi (2017). The amplification factor for the geometric optics regime with the SPA without considering the MSD is given by Eq. (9). For a PM lens there are always two images. Their time delay is related to the (dimensionless) source position $y$ by Takahashi (2017)
@@
Since the GW waveform is amplified by the amplification factor given by Eq. (9), the luminosity distance of the source inferred from the waveform will be an effective distance defined by $d_{L,j}^{\rm eff}\equiv d_{L}/\sqrt{\mu_{j}}$ from each image. By taking the ratio of $d_{L,+}^{\rm eff}/d_{L,-}^{\rm eff}$, the magnification ratio between lensed images $\mu_{r}\equiv\mu_{-}/\mu_{+}$ can be obtained, and then the source position $y$ can be evaluated by Sereno et al. (2011)
@@
On the other hand, the amplification factor for the wave-optics regime is computed exactly with Eq. (28) with different expression of $\Psi(x)$ for different lens models. For the PM model Tambalo et al. (2023), the analytical solution for the amplification factor with MSD is found to be Cremonese et al. (2021)
@@
where $\Gamma$ is the gamma function and $\prescript{}{1}{F}_{1}$ is the confluent hypergeometric function. By taking $\lambda=1$, Eq. (33) reduces to the form given by Nakamura and Deguchi (1999); Matsunaga and Yamamoto (2006). Note that this expression also follows the general formula of Eq. (29). Since the amplification factor leads to non-trivial distortions of the waveform, studying the invariance under transformations in $\lambda$ requires evaluating how the waveform changes. We will study this in detail in Sec. IV.
The amplification factor for the PM model can also be derived in terms of Laguerre polynomials $L_{n}(z)$ Ezquiaga et al. (2020)
@@
where $\nu=-iw/2$. This is equivalent to the more common representation in terms of hypergeometric functions such as in Eq. (33), except for the $e^{2\nu\log\theta_{E}}$ term. This term just redefines the reference time Ezquiaga et al. (2020) and it is typically neglected. When we include $\lambda$, we obtain
@@ III.4 Singular Isothermal Sphere
In addition, we also consider the Singular Isothermal Sphere (SIS) model, which represents a first order approximation to the density profile of a galaxy and its dark matter halo (see Binney and Tremaine (1987)). This model is defined by a Newtonian potential $\Psi(x)=x$. The density profile of the SIS model is related to the velocity dispersion $\sigma_{v}$ by $\rho(r)=\sigma_{v}^{2}/2\pi Gr^{2}$. Then the Einstein radius of the SIS model is obtained by Takahashi (2017)
@@
and the redshifted lens mass in the SIS model is Takahashi (2017)
@@
As a result, one can define the dimensionless frequency for the SIS model $w$ similarly to the PM model, which is $w=4\pi GM_{L,z}\omega/c^{3}$. For a SIS model, two images form when the source position is within the Einstein radius, $\theta_{S}<\theta_{*}$. Otherwise, there is only one image. In the multiple-image region, the time delays and the magnification factors of the images in the geometric optics limit are given by Takahashi (2017)
@@
In the wave-optics limit, by inserting the potential $\Psi(x)=x$ in Eq. (25), an analytical solution (in terms of a series expansion) can be obtained as shown by Eq. (32) in Matsunaga and Yamamoto (2006). Under the MSD, we insert $\Psi(x)=x$ into Eq. (28). By observing that the only change in the integral is replacing $w$ by $w\lambda$, the solution of Eq. (28) for the SIS model can be transformed from Eq. (32) in Matsunaga and Yamamoto (2006) as

### v2

@@ I Introduction
Gravitational lensing of gravitational waves (GWs) occurs when the signal passes close to a massive object that deflects it. This effect could result in multiple images of the same event, magnifications, interference patterns, and delays w.r.t. the unlensed signal. With the advancement of the fourth observing run (O4) of the network of LIGO (with the Livingston and Hanford detectors [1]), Virgo [2], and KAGRA [3, 4, 5] (LVK) and in future runs [6], the number of GW detections will grow further, and with it the probability of detecting a lensed event. For this reason, in upcoming observing runs, it will be crucial to be able to account for the potential effects of lensing on observed signals. Neglecting it could introduce biases into our findings regarding the properties of sources [7, 8, 9], as lensing magnification, for example, alters the inferred luminosity distance. Furthermore, under strong lensing conditions, distortions in the lensed signal may occur, particularly when higher modes, precession, or eccentricity are significant factors [10]. In the wave-optics regime (i.e., when the signal size is comparable to the size of the lens), the modulation of the signal could mimic some feature of the source, like, for example, the eccentricity, as well as biasing other parameters [11]. To ensure accurate analyses, it will be necessary to incorporate lensed waveforms into the waveform models for parameter estimation [12, 13, 14, 15, 16]. Alternatively, employing different methods such as Bayesian analysis [17, 18, 19] or deep learning approaches [20, 21, 22] may prove beneficial. Lensing effects also warrant consideration in population studies, as they influence inferred redshift and mass distributions by amplifying signals and allowing the detection of events from greater distances [23, 24].
@@
The first observation of a lensed GW is still elusive after searches in the O1 and O2 runs [13, 25, 26, 16, 27], O3a [28], the full O3 run [29], and an extended O3 follow-up analysis [30] yielded no confident signatures. Nonetheless, there are good reasons to be optimistic about future detections, as many different studies forecast $\sim 1$ strong lensing detection per year [31, 32, 24, 33].
@@
According to [34], the probability of microlensing (i.e. in the wave-optics regime) is $<20\%$ of all detectable events above the threshold signal-to-noise ratio (SNR)111 In the paper, the threshold SNR is set at $\rho=10$. The level of the detectability of a lensed event is influenced by the mismatch between the lensed signal and the (best matching) unlensed waveform because a mismatch $>0$ decreases the detected SNR. for $M_{det}=120\leavevmode\nobreak\ M_{\odot}$ and $<5\%$ for more common events with $M_{det}=60\leavevmode\nobreak\ M_{\odot}$ (where $M_{det}$ is the detector-frame mass).
@@
The detection and subsequent study of lensed GWs is a powerful tool for both astrophysical and cosmological studies. For example, gravitationally lensed events can be useful because (i) contrarily to unlensed GWs, they give an improved localisation of the source [35, 36]; (ii) the magnification due to the lensing effect extends the range of detectability of GWs to higher redshift (due to magnification) (see e.g. [24]); (iii) through the characterisation of the lens, one can study the content and composition of the universe [14, 37, 38, 39]; the study of one or several lensed systems can be exploited for (iv) precision cosmology [40, 41, 42, 26, 35], (v) statistical cosmology [24], and (vi) tests of general relativity [43, 44, 45, 46, 10, 47].
@@
The time delay between multiple images in gravitational lensing has been a long promise to measure the Hubble constant $H_{0}$ [48], since the time-delay distance, which is a combination of the angular diameter distances between source, lens, and observer, is proportional to it. This method, also called the time-delay cosmography, was first proposed by Refsdal in 1964 [49] and can be similarly performed with strongly-lensed GWs. For multiple lensed GW signals from compact binary coalescence, the lens model can be inferred from the magnification ratio and the arrival time delay between them, using ground-based detectors [41, 35, 50], space-based antennas [40], or pulsar timing arrays [51]. If accompanied by electromagnetic (EM) counterparts, the host galaxies of LVK GW sources can be pinpointed using the information from galaxy catalogs [41, 35], so that $H_{0}$ could be measured at high accuracy.
@@
Unfortunately, the usefulness of studying lensed events has a large caveat that goes under the name of mass-sheet degeneracy (MSD) [52, 53, 54, 55]. The MSD relies on the fact that the simultaneous scaling of the lens mass and the source plane leaves the geometrical lensing outputs unchanged. Thus, in practice, we have an intrinsic indeterminacy when we analyze observational data to reconstruct the mass (model) of the lens from the image structure. This is a well-known problem in electromagnetic (EM) lensing.
@@
In EM studies, the MSD is always defined in the geometrical optics approximation, when the wavelength of the lensed signal is much smaller than the typical size of the lens. In this limit, the MSD can be broken, or at least attenuated, when: (i) the same gravitational object act as lens for multiple sources at different redshifts [56, 57, 58, 41]; (ii) independent mass estimations of the lens are available, e.g. from spatially resolved kinematic observations of the lens [54, 59, 60]; (iii) by a proper combination of the previous astrophysical data with cosmological geometrical data [61]. Further discussions in strong-lensing degeneracies can also be found in [62, 63, 64].
@@
In the context of gravitational lensing of GWs, given the much lower frequency of the signal, the wave optics regime can also be relevant as GW wavelengths can be comparable to the lens size [65, 66, 51, 67, 68, 69, 70]. Even though the MSD can be broken in this limit [71], such a degeneracy can affect lens mass modeling and can highly bias the estimation of the Hubble constant, $H_{0}$ [72, 54, 73], contributing to the systematic error budget. This is the reason why the MSD is carefully analyzed in projects like H0LiCOW [74, 75], TDCOSMO [76] and in the Time Delay Lens modeling Challenge [77, 78]. In GW lensing, this topic was first discussed in [71] and, more recently in the geometric optics regime in [79].
@@ II.1 General phenomenology and formalism
Therefore, under these approximations, the problem simplifies to that of solving a scalar wave equation on a medium, i.e. a Helmholtz equation (see e.g. [80] for a review):
@@
Apart from the phase shift of type II images that can induce waveform distortions on signals with multiple frequency components [10], lensing is achromatic in the SPA.
@@
In other words, weak-field lensing occurs at scales much larger than the strong gravity regime unless the source is very close to the lens, $D_{LS}\ll D_{L},D_{S}$, as is the case e.g. for the photons near the super-massive black holes imaged by the Event Horizon Telescope [81, 82]. For typical astrophysical lensing situations in which $D_{LS}\approx D_{L}\approx D_{S}$, then
@@ II.2 GW lensing
  • the signal is understood from first principles: general relativity can be used to derive waveform models for CBCs (though numerical simulations and various effective modeling approaches are needed in practice, see [83]),
@@
All these characteristics will be relevant for different aspects of gravitational lensing. Importantly, GW sky localizations, which are at best of the order of few square degrees in optimistic current observing scenarios [6], will be much larger than the typical angular scale of the problem $\theta_{*}$.
@@ III Invariance Transformations in gravitational wave lensing
Our ability to learn about astrophysics and cosmology from lensed systems depends on how well we can distinguish the predictions of different models on the observables. For this reason, we now focus on studying which sets of variations of the lens system leave the observables unchanged, as those will lead to degeneracies between different astrophysical and cosmological interpretations. Such “invariance transformations” have been widely studied in the EM spectrum since they constitute a large fraction of the error budget when reconstructing the lens model or inferring the expansion rate of the Universe [52, 53, 54, 55]. They have also been explored in the context of GWs [71, 79], and here we will build upon those results.
@@ III.1 Multiple-image regime
Following [54], we are going to consider three possible modifications of our fiducial lens scenario:
@@ III.2 Wave-optics regime
We now explore these invariance transformations in the regime of wave optics. Since the diffraction integral is not analytically solvable in general, we restrict to the simpler case of spherically symmetric lenses. In this case, the diffraction integral in Eq. (7) can be simplified to one dimension, integrating over the angular polar coordinate. Doing so, one arrives at [80]
@@ III.3 Point mass model
To gain further intuition, we fully specify the lens model to exemplify the effect of the invariance transformations. We start with the simplest model: a point mass (PM), where the potential is given by $\Psi(x)=\log x$. The Einstein radius in the PM model is $\theta_{*}=\sqrt{4GM_{l,z}/c^{3}\tau_{D}}$ [65]. The amplification factor for the geometric optics regime with the SPA without considering the MSD is given by Eq. (9). For a PM lens there are always two images. Their time delay is related to the (dimensionless) source position $y$ by [65]
@@
Since the GW waveform is amplified by the amplification factor given by Eq. (9), the luminosity distance of the source inferred from the waveform will be an effective distance defined by $d_{L,j}^{\rm eff}\equiv d_{L}/\sqrt{\mu_{j}}$ from each image. By taking the ratio of $d_{L,+}^{\rm eff}/d_{L,-}^{\rm eff}$, the magnification ratio between lensed images $\mu_{r}\equiv\mu_{-}/\mu_{+}$ can be obtained, and then the source position $y$ can be evaluated by [40]
@@
On the other hand, the amplification factor for the wave-optics regime is computed exactly with Eq. (28) with different expression of $\Psi(x)$ for different lens models. For the PM model [84], the analytical solution for the amplification factor with MSD is found to be [71]
@@
where $\Gamma$ is the gamma function and $\prescript{}{1}{F}_{1}$ is the confluent hypergeometric function. By taking $\lambda=1$, Eq. (33) reduces to the form given by [80, 85]. Note that this expression also follows the general formula of Eq. (29). Since the amplification factor leads to non-trivial distortions of the waveform, studying the invariance under transformations in $\lambda$ requires evaluating how the waveform changes. We will study this in detail in Sec. IV.
The amplification factor for the PM model can also be derived in terms of Laguerre polynomials $L_{n}(z)$ [86]
@@
where $\nu=-iw/2$. This is equivalent to the more common representation in terms of hypergeometric functions such as in Eq. (33), except for the $e^{2\nu\log\theta_{E}}$ term. This term just redefines the reference time [86] and it is typically neglected. When we include $\lambda$, we obtain
@@ III.4 Singular Isothermal Sphere
In addition, we also consider the Singular Isothermal Sphere (SIS) model, which represents a first order approximation to the density profile of a galaxy and its dark matter halo (see [87]). This model is defined by a Newtonian potential $\Psi(x)=x$. The density profile of the SIS model is related to the velocity dispersion $\sigma_{v}$ by $\rho(r)=\sigma_{v}^{2}/2\pi Gr^{2}$. Then the Einstein radius of the SIS model is obtained by [65]
@@
and the redshifted lens mass in the SIS model is [65]
@@
As a result, one can define the dimensionless frequency for the SIS model $w$ similarly to the PM model, which is $w=4\pi GM_{L,z}\omega/c^{3}$. For a SIS model, two images form when the source position is within the Einstein radius, $\theta_{S}<\theta_{*}$. Otherwise, there is only one image. In the multiple-image region, the time delays and the magnification factors of the images in the geometric optics limit are given by [65]
@@
In the wave-optics limit, by inserting the potential $\Psi(x)=x$ in Eq. (25), an analytical solution (in terms of a series expansion) can be obtained as shown by Eq. (32) in [85]. Under the MSD, we insert $\Psi(x)=x$ into Eq. (28). By observing that the only change in the integral is replacing $w$ by $w\lambda$, the solution of Eq. (28) for the SIS model can be transformed from Eq. (32) in [85] as